Ask your own question, for FREE!
Mathematics 13 Online
OpenStudy (anonymous):

Assume that there are 6 different issues of Sports Illustrated, 7 different issues of Time, and 3 different issues of Newsweek, including the December 1st issue, on the rack. You choose 4 of them at random. (1) What is the probability that exactly 1 is an issue issue of Sports Illustrated? (2) What is the probability that you choose the December 1st issue of Newsweek?

Directrix (directrix):

(1) [C(6,1) C(10,3)] / C(16,4) = 36/91 = .396 approx

Directrix (directrix):

(2) [C(1,1) C(15,3)] / C(16,4) = 1/4 = .25

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
3 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
3 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!